OpenStudy (anonymous):

if ten students from a class have volunteered to organize a dance, in how many ways can six be chosen for the committee?

OpenStudy (anonymous):

you are being asked for "ten choose six" written as \(\dbinom{10}{6}\) and computed via \[\dbinom{10}{6}=\frac{10\times 9\times 8\times 7}{4\times 3\times 2}\] cancel first, multiply last